When discharging the 20.0 mF capacitor, you measur
Answer #1

7.

Vr - Vc = 0

Vr = Vc = 11.2 V

8.
I = Vr / R = 11.2 / 300 =0.037 A


9. Vc = V [ e^(-t/RC) ]

so this is Exponential decay.